# MAC-VO: Metrics-aware Covariance for Learning-based Stereo Visual Odometry mac-vo.github.io

### v1

@@ Abstract
We propose the MAC-VO, a novel learning-based stereo VO that leverages the learned metrics-aware matching uncertainty for dual purposes: selecting keypoint and weighing the residual in pose graph optimization. Compared to traditional geometric methods prioritizing texture-affluent features like edges, our keypoint selector employs the learned uncertainty to filter out the low-quality features based on global inconsistency. In contrast to the learning-based algorithms that model the scale-agnostic diagonal weight matrix for covariance, we design a metrics-aware covariance model to capture the spatial error during keypoint registration and the correlations between different axes. Integrating this covariance model into pose graph optimization enhances the robustness and reliability of pose estimation, particularly in challenging environments with varying illumination, feature density, and motion patterns. On public benchmark datasets, MAC-VO outperforms existing VO algorithms and even some SLAM algorithms in challenging environments. The covariance map also provides valuable information about the reliability of the estimated poses, which can benefit decision-making for autonomous systems.
@@ I Introduction
Visual Odometry (VO) predicts the relative camera pose from image sequences and often serves as the front-end of Simultaneous Localization and Mapping (SLAM) systems. Over the past few decades, both geometric and learning-based methods have been developed with significant advances in generalizability and accuracy [1, 2, 3, 4]. However, VO remains a challenging problem in real-world scenarios, with multiple visual degraded scenarios such as low illumination, dynamic and texture-less scenes.
To improve the robustness in challenging scenes, geometric-based VO algorithms employ outlier filtering strategies [5] and weigh the optimization residuals by the covariance of the observed features [6]. However, how to effectively select the reliable keypoints and model their covariance becomes two significant challenges. Existing methods typically select the keypoints based on local intensity gradient with a manually defined threshold [7, 8, 9]. These approaches leads to errors and outliers because it doesn’t model the structure or context information of the environment (e.g. features on repetitive patterns may not be ideal candidates despite high image gradients). Moreover, the covariance model is often empirically modeled using a constant parameter, which is sub-optimal. In addition, the parameters in keypoint selection and covariance model need to be extensively tuned for different environments.
Figure 1: Without relying on multi-frame bundle adjustment, MAC-VO aggregately reconstruct the map based on the two-frame Pose graph optimization. We propose metrics-aware covariance model for 3D keypoint based on the learning matching uncertainty. The covariance model serves dual purposes: selecting keypoints and weighing the residual in the pose graph optimization.
@@
  • We present a learning-based 2D uncertainty network with metrics awareness, leveraging iterative motion aggregator to capture the inconsistency of the feature matching. This metrics-aware uncertainty evaluates the quality of features in the keypoint selector and weights the residual of backend optimization.
@@
Figure 2: MAC-VO pipeline. We use a shared matching network to estimate the depth, flow, and corresponding uncertainty. The learned uncertainty is leveraged to filter out unreliable features and model the metrics-aware covariance of the 3D keypoints. These registered keypoints and their covariance models are then utilized in the back-end optimization to determine the relative camera pose.
@@ II Related Works
Existing geometric-based methods optimize the camera pose based on geometric constraints like re-projection error [1, 13] or photometric error [14, 15, 16]. To more accurately model the uncertainty of the depth, Civera et al. [17] and Montiel[18] investigate the inverse depth parameterization. These methods often use constant parameters and simple heuristics to model the covariance matrix of these errors during the factor graph optimization. For multi-sensor fusion [6, 19] and semantic SLAM [20, 21], the covariance model plays a significant role in weighing the confidence of different sensors and modules. To effectively capture sensor uncertainty, the covariance models are often tuned based on empirical prior. However, these simplified covariance models fail to capture the complexity of the challenging environments.
@@
Amidst these developments, hybrid learning-based SLAM systems [32, 26, 33, 34, 35, 36] are emerging to synergize the geometric constraint with the adaptability of learning-based methods. To improve the reliability of the feature tracking process, some methods introduce learning-based uncertainty measurements or confidence scores [37, 38, 10] for pose optimization. DROID-SLAM [3] utilizes a differentiable bundle adjustment layer to implicitly tune the uncertainty model. This method employs a simplified diagonal covariance model for the bundle adjustment. These methods focus on the relative confidence between features , but ignore the scale consistency of the covariance model.
@@ III Method
As illustrated in Figure 2, MAC-VO outlines an effective integration of a learning-based front-end and a geometrically constrained back-end using the metrics-aware covariance model. In the front-end (Section III-A), we train an uncertainty-aware matching network to model the corresponding uncertainty stems from feature deficiencies. Utilizing the learned uncertainty, we develop a keypoint selector in Section III-B to choose reliable features. The metrics-aware 2D uncertainty is then propagated to the 3D space via the proposed covariance model in Section. III-C. In the back-end optimization (Section. III-D), we optimize the relative motion by minimizing the distance between registered keypoints weighted by the 3D covariance.
Figure 4: Top: Architecture of the uncertainty-aware matching network. We employ a motion aggregator and an iterative update structure to enable the covariance module to capture inconsistencies in feature matching. Bottom: In each iteration, the model captures the inconsistency between the motions. For the $\Delta\sigma$, the red color indicates a positive $\Delta\sigma$ that increases the uncertainty, and blue means decreasing the uncertainty.
@@ III-A Network & Uncertainty Training
The objective of our network is to predict the flow $\hat{f}\in\mathbb{R}^{2}$ and the corresponding uncertainty $\hat{\Sigma}_{f}=\mathrm{diag}(\sigma_{u}^{2},\sigma_{v}^{2})$. As shown in Fig. 4, an iterative motion aggregator inspired by the FlowFormer [11] is employed to capture the inconsistency in feature matching. To extend this network for uncertainty estimation, we add an uncertainty decoder to predict the $\Delta\sigma$ updates of the uncertainty in the log space. The use of log space enables additive updates and constrains the output range, stabilizing gradients and simplifying model output. After iterative updates, the log-variance passes through an exp activation function to obtain the final uncertainty. More details about the network are shown in Appendix. B-E.
@@
$\mathcal{L}_{cov}=\sum_{i}^{N}\alpha_{i}\left((y-f_{i})^{\top}\hat{\Sigma}_{f}^{-1}(y-f_{i})+\log(\det\hat{\Sigma}_{f})\right),$ (1)
where $y$ is the ground truth optical flow, $f_{i}$ denotes the $i$-th iteration of the network outputs, and $\alpha_{i}$ is the weight for each iteration and is set to decrease exponentially with ratio of $0.8$. During the training stage, we initialize the encoder network parameters with the pre-trained model by FlowFormer. We then train the covariance module on the synthetic dataset TartanAir [43]. Our experiments demonstrate that the model can generalize to real-world datasets without fine-tuning.
@@ III-B Uncertainty-based Keypoint Selection
Different from the random selector used in DPVO[4] and the hand-crafted features used in ORB-SLAM [8], we leverage the learned uncertainty estimation to filter out unreliable features such as those on the vehicle illustrated in Fig. 3. This is achieved by composing three filters: the uncertainty filter, geometry filter, and the non-minimum suppression (NMS). The uncertainty filter removes pixels with depth and flow uncertainty larger than $1.5$ times the median uncertainty of the current frame, which discards the unreliable features while maintaining the diversity of keypoint candidates. The accurate uncertainty estimation effectively removes the keypoints on occluded objects, reflective surfaces, and feature-less areas. Illustrated in Fig. 3, the uncertainty filter removes all keypoint candidates on the moving vehicle on a KITTI trajectory due to its high flow uncertainty. Along with the uncertainty filter is the geometric filter, which constraints the disparity and removes keypoints on the edge of frame. To ensure the even distribution of keypoints in image, the NMS filter is applied on the product of depth and flow uncertainty map prior to both filters.
@@ III-C Metrics-aware 3D Covariance Model
In the context of the camera projection geometry, the covariance of a 3D keypoint is determined by the uncertainty of the depth $\sigma_{d}^{2}$ and matching $(\sigma_{u}^{2},\sigma_{v}^{2})$. To accurately model the covariance for 3D keypoint, it is critical to determine (1) the depth uncertainty of the matched points $\hat{\sigma}_{d}^{2}$ and (2) the off-diagonal covariance terms during the 2D-3D projection process.
Depth Uncertainty after keypoint matching  As shown in Fig. 5 (a), the matched features are expected to be within a probabilistic range centered at $[u_{i,t},v_{i,t}]$ due to flow uncertainty $(\sigma_{u_{i,t}}^{2},\sigma_{v_{i,t}}^{2})$. As a result, a minor disturbance in the feature matching may introduce a large error in the depth.
To address this problem, we model the depth uncertainty $\hat{\sigma}_{d_{i,t}}^{2}$ of the matched feature point based on the depth feature of the local patch $D_{i,t}$. We approximate it with the weighted sum of the variances within the patch. The weights are determined by a 2D Gaussian kernel $\varphi$, which utilizes $\sigma_{u_{i,t}}^{2}$ and $\sigma_{v_{i,t}}^{2}$ to adjust the influence of each point within the patch: $\sigma_{d_{i,t}}^{2}=\sum_{j}{\varphi_{j}(d_{j}-\mu_{{D}_{i,t}})^{2}}.$
Project 2D Uncertainty to 3D Covariance  Following the pinhole camera model with focus $\mathrm{f}_{x},\mathrm{f}_{y}$ and optical center $\mathrm{c}_{x},\mathrm{c}_{y}$, the coordinate of the keypoint is calculated by: $\mathbf{x}_{i,t}=(\mathbf{u}_{i,t}-\mathrm{c}_{x})\mathbf{d}_{i,t}/\mathrm{f}_{x}$, $\mathbf{y}_{i,t}=(\mathbf{v}_{i,t}-\mathrm{c}_{y})\mathbf{d}_{i,t}/\mathrm{f}_{y}$ and $\mathbf{z}_{i,t}=\mathbf{d}_{i,t}$, as shown in Fig. 5 (b). To accurately capture the uncertainties associated with these measurements, the main diagonal of the covariance matrix is formulated as:
$\sigma_{x_{i,t}}^{2}$ $={((\sigma_{u_{i,t}}^{2}+d^{2}_{i,t})(\sigma_{d_{i,t}}^{2}+u^{2}_{i,t})-u^{2}_{i,t}d^{2}_{i,t}+\mathrm{c}_{x}^{2}\sigma_{d_{i,t}}^{2})}/{\mathrm{f}_{x}^{2}},$ (2)
$\sigma_{y_{i,t}}^{2}$ $={((\sigma_{v_{i,t}}^{2}+d_{i,t}^{2})(\sigma_{d_{i,t}}^{2}+v_{i,t}^{2})-v_{i,t}^{2}d_{i,t}^{2}+\mathrm{c}_{y}^{2}\sigma_{d_{i,t}}^{2})}/{\mathrm{f}_{y}^{2}},$
@@
${}^{c}\Sigma_{i,t}^{p}=\begin{bmatrix}\sigma_{z}^{2}&\sigma_{xz_{i,t}}&\sigma_{yz_{i,t}}\\ \sigma_{xz_{i,t}}&\sigma_{x_{i,t}}^{2}&\sigma_{xy_{i,t}}\\ \sigma_{yz_{i,t}}&\sigma_{xy_{i,t}}&\sigma_{y_{i,t}}^{2}\\ \end{bmatrix},\;\begin{aligned} &\sigma_{xz}=\sigma_{d}^{2}(u-\mathrm{c}_{x})/\mathrm{f}_{x}\\ &\sigma_{yz}=\sigma_{d}^{2}(v-\mathrm{c}_{y})/\mathrm{f}_{y}\\ &\sigma_{xy}=\frac{\sigma_{d}^{2}(u-\mathrm{c}_{x})(v-\mathrm{c}_{y})}{\mathrm{f}_{x}\mathrm{f}_{y}}.\end{aligned}$ (3)
Results in the ablation study also confirm the necessity of off-diagonal terms for accurate pose graph optimization. Detailed derivation for Eq. 2 and Eq. 3 is in Appendix. A-C.
@@ III-D Pose Graph Optimization
$T^{\star}$ $=\operatorname*{arg\,min}_{T_{t}}{\sum_{i}{\left\|p_{i,t-1}-{T}_{t}\,{}^{c}p_{i,t}\right\|^{2}_{\Sigma_{i}}}},$ (4)
@@
$\|\cdot\|_{\Sigma_{i}}$ represents the Mahalanobis distance with covariance matrix $\Sigma_{i}$. Unlike DROID-SLAM [3], which employs a diagonal covariance matrix, we model the correlation between axes to capture accurate inter-dependencies. We solved this optimization problem by Levenberg-Marquardt algorithms using PyPose [44].
@@ IV Experiment
Datasets & Baseline  We evaluate the proposed model and baseline methods on a variety of public datasets, including synthetic dataset TartanAir v2 [43], real-world data from EuRoC [45], KITTI [46], as well as customized data collected from a Zed camera. These datasets cover a diverse range of hardware configurations, motion patterns, and environments. To demonstrate our method’s robustness under challenging scenarios, we collected the TartanAir v2, a new set of difficult trajectories following the TartanAir [43] that includes frequent indoor-outdoor transition and low-illumination scenes as shown in Fig. 6. To demonstrate the generalizability of our model, we use the same configuration across all datasets.
Evaluation Metrics  Our evaluation focuses more on relative error Since the proposed method does not contain loop closure or global bundle adjustment. So we use Relative translation error ($t_{\mathrm{rel}}$, $\mathrm{m/frame}$) and relative rotation error ($r_{\mathrm{rel}}$, $\mathrm{{}^{\circ}/frame}$) as:
$t_{\mathrm{rel}}$ $=\frac{1}{N}\sum^{N}_{t=1}{\left\|p_{t+1}-p_{t}-R_{t}\hat{R}_{t}^{\top}\!\!\left(\hat{p}_{t+1}-\hat{p_{t}}\right)\right\|_{2}},$ (5)
@@
where $p_{t}$ and $R_{t}$ are ground truth position and rotation, $\hat{p}_{t}$ and $\hat{R}_{t}$ is the estimated position and rotation. $R_{t,t+1}=R_{t}^{\top}R_{t+1}$ is the rotation from frame $t$ to frame $t+1$.

### v2

@@ Abstract
We propose MAC-VO, a novel learning-based stereo visual odometry (VO) framework that trains a metrics-aware uncertainty model to serve two critical functions: selecting keypoints and weighting residuals in pose graph optimization. Unlike traditional geometric methods that favor texture-rich features like edges, our keypoint selector leverages this learned uncertainty model to eliminate low-quality features based on global inconsistency. In contrast to learning-based approaches that rely on scale-agnostic weight matrices for covariance, our metrics-aware covariance model—derived from the learned uncertainty—captures spatial errors in keypoint registration and inter-axis correlations. By embedding this covariance model into pose graph optimization, MAC-VO achieves superior robustness and accuracy in pose estimation, excelling in challenging environments with varying illumination, feature density, and motion patterns. Evaluations on public benchmark datasets demonstrate that MAC-VO surpasses existing VO algorithms and even some SLAM systems in difficult scenarios. Additionally, the uncertainty map offers valuable insights for decision-making.
@@ I Introduction
Visual Odometry (VO) predicts the relative camera pose from image sequences and often serves as the front-end of Simultaneous Localization and Mapping (SLAM) systems. Over the past few decades, both geometric and learning-based methods have been developed with significant advances in generalizability and accuracy [1, 2, 3, 4]. However, VO remains a challenging problem in real-world scenarios, with multiple visually degraded scenarios such as low illumination, dynamic and texture-less scenes.
To improve the robustness in challenging scenes, geometric-based VO algorithms employ outlier filtering strategies [5] and weigh the optimization residuals by the covariance of the observed features [6]. However, how to effectively select the reliable keypoints and model their covariance becomes two significant challenges. Existing methods typically select the keypoints based on local intensity gradient with a manually defined threshold [7, 8, 9]. These approaches lead to errors and outliers because they don’t model the structure or context information of the environment (e.g., features on repetitive patterns may not be ideal candidates despite high image gradients). Moreover, the covariance model is often empirically modeled using a constant parameter, which is sub-optimal. In addition, the parameters in the keypoint selection and covariance model need to be extensively tuned for different environments.
Figure 1: Without relying on multi-frame bundle adjustment, MAC-VO aggregately reconstructs the map based on the two-frame Pose graph optimization. We propose metrics-aware covariance model for 3D keypoint based on the learning matching uncertainty.
@@
  • We present a learning-based 2D uncertainty network with metrics awareness, leveraging an iterative motion aggregator to capture the inconsistency of the feature matching. We use this metrics-aware uncertainty to evaluate the quality of features, in order to select keypoints, and guide the backend optimization.
@@
Figure 2: MAC-VO pipeline. It employs a shared matching network to jointly estimate depth, optical flow, and corresponding uncertainties. The learned uncertainty is utilized in the keypoint selector, the metrics-aware 3D covariance model, and the back-end optimization.
@@ II Related Works
Existing geometric-based methods optimize the camera pose based on geometric constraints like re-projection error [1, 13] or photometric error [14, 15, 16]. To more accurately model the uncertainty of the depth, Civera et al. [17] and Montiel[18] investigate the inverse depth parameterization. These methods often use constant parameters and simple heuristics to model the covariance matrix of these errors during the factor graph optimization. For multi-sensor fusion [6, 19] and semantic SLAM [20, 21], the covariance model plays a significant role in weighting the confidence of different sensors and modules. To capture sensor uncertainty, the covariance models are often tuned based on empirical prior. However, these simplified covariance models fail to capture the complexity of the challenging environments.
@@
Amidst these developments, hybrid learning-based SLAM systems [32, 26, 33, 34, 35, 36] are emerging to synergize the geometric constraint with the adaptability of learning-based methods. To improve the reliability of the feature tracking process, some methods introduce learning-based uncertainty measurements or confidence scores [37, 38, 10, 39, 40] for pose optimization. DROID-SLAM [3] utilizes a differentiable bundle adjustment layer to implicitly tune the uncertainty model. This method employs a simplified diagonal covariance model for the bundle adjustment. These methods focus on the relative confidence between features, but ignore the scale consistency of the covariance model.
@@ III Method
As illustrated in Fig. 2, MAC-VO outlines an effective integration of a learning-based front-end and a geometrically constrained back-end using the metrics-aware covariance model. In the front-end (Section III-A), we train an uncertainty-aware matching network to model the corresponding uncertainty stems from feature deficiencies. Utilizing the learned uncertainty, we develop a keypoint selector in Section III-B to choose reliable features. The metrics-aware 2D uncertainty is then propagated to the 3D space via the proposed covariance model in Section III-C. In the back-end optimization (Section III-D), we optimize the relative motion by minimizing the distance between registered keypoints weighted by the 3D covariance.
Figure 4: Top: Architecture of the uncertainty-aware matching network. Bottom: Each iteration, the model captures the inconsistency of the matching. For the $\Delta\sigma$, the red color indicates a positive $\Delta\sigma$ increasing the uncertainty, and blue means decreasing uncertainty.
@@ III-A Network & Uncertainty Training
The objective of the proposed matching network is to estimate the flow $\hat{f}\in\mathbb{R}^{2}$ and the corresponding uncertainty $\hat{\Sigma}_{f}=\mathrm{diag}(\sigma_{u}^{2},\sigma_{v}^{2})$ between two frames. We adopt the FlowFormer architecture [11] as the backbone for optical flow estimation, leveraging its transformer-based design. As shown in Fig. 4, the recurrent decoder iteratively refines flow predictions by capturing ambiguities in feature correspondence within the cost volume’s matching space, utilizing global context from the encoded cost memory.
To extend this framework and leverages both global motion cues and local feature, we introduce a covariance decoder that predicts $\log\Delta\sigma$, the iterative updates of the uncertainty on log space. Operating the uncertainty update on the log space facilitates additive updates, ensures positive variances, and stabilizes gradients. After iterative updates, the log-variance passes through an exp activation function to obtain the final uncertainty. More details about the network and training are shown in Appendix B-E.
@@
$L_{cov}=\sum_{i}^{N}\alpha_{i}\left((y-\hat{f}_{i})^{\top}\hat{\Sigma}_{f}^{-1}(y-\hat{f}_{i})+\log(\det\hat{\Sigma}_{f})\right),$ (1)
where $y$ is the ground truth optical flow, $f_{i}$ denotes the $i$-th iteration of the network outputs, and $\alpha_{i}$ is the weight for each iteration and is set to decrease exponentially with a ratio of $0.8$. During the training stage, we initialize the encoder network parameters with the pre-trained model by FlowFormer. We then train the covariance module on the synthetic dataset TartanAir [45], and evaluate the model in the TartanAirv2 sequences and other real-world datasets. Our experiments demonstrate that the model can generalize to real-world datasets without fine-tuning.
@@ III-B Uncertainty-based Keypoint Selection
Different from the random selector used in DPVO[4] and the hand-crafted features selector used in ORB-SLAM [8], we leverage the learned uncertainty estimation to filter out unreliable features. This is achieved by composing three filters: the uncertainty filter, geometry filter, and the non-minimum suppression (NMS). To prevent the clustering of keypoint candidates, we first apply the NMS filter, ensuring a more even spatial distribution of features. The geometry filter further removes keypoints located at the image boundaries and those outside the valid depth observation range. Finally, the uncertainty filter eliminates pixels with depth and flow uncertainty greater than 1.5 times the median uncertainty of the current frame. Illustrated in Fig. 3, the uncertainty filter removes all keypoint candidates on the moving vehicle on a KITTI trajectory. Our proposed uncertainty estimation strategy also filters out keypoints on occluded objects, reflective surfaces, and feature-less areas, resulting in a more robust and reliable feature set.
@@ III-C Metrics-aware 3D Covariance Model
In the context of the camera projection geometry, the covariance of a 3D keypoint is determined by the uncertainty of the depth $\sigma_{d}^{2}$ and matching $(\sigma_{u}^{2},\sigma_{v}^{2})$. To accurately model the covariance for 3D keypoint, it is critical to determine (1) the depth uncertainty of the matched points $\hat{\sigma}_{d}^{2}$ and (2) the off-diagonal covariance terms during the 2D-3D projection.
Project 2D Uncertainty to 3D Covariance  Following the pinhole camera model with focal length $\mathrm{f}_{x},\mathrm{f}_{y}$ and optical center $\mathrm{c}_{x},\mathrm{c}_{y}$, the coordinate of the keypoint is calculated by: $\mathbf{x}_{i,t}=(\mathbf{u}_{i,t}-\mathrm{c}_{x})\mathbf{d}_{i,t}/\mathrm{f}_{x}$, $\mathbf{y}_{i,t}=(\mathbf{v}_{i,t}-\mathrm{c}_{y})\mathbf{d}_{i,t}/\mathrm{f}_{y}$ and $\mathbf{z}_{i,t}=\mathbf{d}_{i,t}$, as shown in Fig. 5 (a). To accurately capture the uncertainties associated with these measurements, the main diagonal of the covariance matrix is formulated as:
$\sigma_{x_{i,t}}^{2}$ $={\left(\sigma_{u_{i,t}}^{2}\sigma_{d_{i,t}}^{2}+\sigma_{u_{i,t}}^{2}d_{i,t}^{2}+(u_{i,t}-c_{x})^{2}\sigma_{d_{i,t}}^{2}\right)}/{\mathrm{f}_{x}^{2}},$ (2)
$\sigma_{y_{i,t}}^{2}$ $={\left(\sigma_{v_{i,t}}^{2}\sigma_{d_{i,t}}^{2}+\sigma_{v_{i,t}}^{2}d_{i,t}^{2}+(v_{i,t}-c_{y})^{2}\sigma_{d_{i,t}}^{2}\right)}/{\mathrm{f}_{y}^{2}},$
@@
${}^{c}\Sigma_{i,t}^{p}=\begin{bmatrix}\sigma_{z}^{2}&\sigma_{xz_{i,t}}&\sigma_{yz_{i,t}}\\ \sigma_{xz_{i,t}}&\sigma_{x_{i,t}}^{2}&\sigma_{xy_{i,t}}\\ \sigma_{yz_{i,t}}&\sigma_{xy_{i,t}}&\sigma_{y_{i,t}}^{2}\\ \end{bmatrix}\begin{aligned} &\sigma_{xz}=\sigma_{d}^{2}(u-\mathrm{c}_{x})/\mathrm{f}_{x}\\ &\sigma_{yz}=\sigma_{d}^{2}(v-\mathrm{c}_{y})/\mathrm{f}_{y}\\ &\sigma_{xy}=\frac{\sigma_{d}^{2}(u-\mathrm{c}_{x})(v-\mathrm{c}_{y})}{\mathrm{f}_{x}\mathrm{f}_{y}}.\end{aligned}$ (3)
Results in the ablation study also confirm the necessity of off-diagonal terms for accurate pose graph optimization. Detailed derivation for Eq. 2 and Eq. 3 is in Appendix A-C.
Uncertainty Correction after Keypoint Matching  As shown in Fig. 5 (b), the matched features are expected to fall in a probabilistic range centered at $[u_{i,t},v_{i,t}]$ with the flow uncertainty $(\sigma_{u_{i,t}}^{2},\sigma_{v_{i,t}}^{2})$. However, the depth observations of matched keypoints may vary due to scene geometry. For example, if a matched keypoint is located near the edge of an object, such as a wall, even a minor perturbation in feature matching can lead to substantial discrepancies in 3D keypoint registration. This, in turn, amplifies the uncertainty in the reconstructed 3D structure.
To address this problem, we correct the depth uncertainty $\hat{\sigma}_{d_{i,t}}^{2}$ of the matched feature point based on the depth feature of the local patch $D_{i,t}$ with the kernel size of $32$. We approximate it with the weighted sum of the variances within the patch. The weights are determined by a 2D Gaussian kernel $\varphi$, which utilizes $\sigma_{u_{i,t}}^{2}$ and $\sigma_{v_{i,t}}^{2}$ to adjust the influence of each point within the patch: $\sigma_{d_{i,t}}^{2}=\sum_{j}{\varphi_{j}(d_{j}-\mu_{{D}_{i,t}})^{2}}.$
@@ III-D Pose Graph Optimization
$T^{\star}$ $=\operatorname*{arg\,min}_{T_{t}}{\sum_{i}{\left\|\mathbf{p}_{i,t-1}-{T}_{t}\,{}^{c}\mathbf{p}_{i,t}\right\|^{2}_{\Sigma_{i}}}},$ (4)
@@
$\|\cdot\|_{\Sigma_{i}}$ represents the Mahalanobis distance with covariance matrix $\Sigma_{i}$. Unlike DROID-SLAM [3], which employs a diagonal covariance matrix, we model the correlation between axes to capture accurate inter-dependencies. We solved this problem by Levenberg-Marquardt using PyPose [46, 47].
@@ IV Experiment
Datasets & Baseline  We evaluate the proposed model and baseline methods on a variety of public datasets, including synthetic dataset TartanAir v2 [45], real-world data from EuRoC [48], KITTI [49], as well as customized data collected from a Zed-X camera. These datasets cover a diverse range of hardware configurations, motion patterns, and environments. To demonstrate our method’s robustness under challenging scenarios, we collected the TartanAir v2, a new set of difficult trajectories following the TartanAir [45] that includes frequent indoor-outdoor transition and low-illumination scenes as shown in Fig. 6. To demonstrate the generalizability of our model, we use the same configuration and same matching model across all datasets.
Evaluation Metrics  Since the proposed method does not contain loop closure or global bundle adjustment, our evaluation focuses on relative error. We use relative translation error ($t_{\mathrm{rel}}$, $\mathrm{m/frame}$) and rotation error ($r_{\mathrm{rel}}$, $\mathrm{{}^{\circ}/frame}$) as:
$t_{\mathrm{rel}}$ $=\frac{1}{N}\sum^{N}_{t=1}{\left\|\mathbf{p}_{t+1}-\mathbf{p}_{t}-R_{t}\hat{R}_{t}^{\top}\!\!\left(\hat{\mathbf{p}}_{t+1}-\hat{\mathbf{p}_{t}}\right)\right\|_{2}},$ (5)
@@
where $\mathbf{p}_{t}$ and $R_{t}$ are ground truth position and rotation, $\hat{\mathbf{p}}_{t}$ and $\hat{R}_{t}$ is the estimated position and rotation. $R_{t,t+1}=R_{t}^{\top}R_{t+1}$ is the rotation from frame $t$ to frame $t+1$.